A student selects one even number and one odd number from
the first 10 even numbers and first 11 odd numbers. How
many combinations can there be?
1
Expert's answer
2020-11-12T18:39:31-0500

Imagine, that student choose on even number. It means he can choose 1 odd number from 11 or 11 different combination. If stufent another even he alse has 11 different combination. He has 10 different even number. It means he has "\underbrace{11+11+...+11}_{10} = 10 * 11 = 110" combinations are there.